NicktheBlade16:

Algebra 2: Compare the graphs below of the logarithmic functions. Write the equation to represent g(x). a graph of f of x equals log x and a graph of a similar log function, g of x, that is the same as f of x and shifted down 4 units g(x) = log(x) − 4 g(x) = log(x) + 4 g(x) = log(x + 4) g(x) = log(x − 4)

NicktheBlade16:

@Alexis1415

Alexis1415:

Ok so again The answer is B, g(x) = log(x) + 4 Reasoning: \[g(x)=f(x)+4\] \[g(x)=\log(x)+4\] So 1 is usually what x is so i'm going to put that into the equation. \[f(1)=\log(1)=0\] \[g(1)=\log(1)+4=4\] So thats how we get, g(x) = log(x) + 4

NicktheBlade16:

Alright. Thank you, i hope you have your question answered and that you have a good day.

Alexis1415:

@nicktheblade16 wrote:
Alright. Thank you, i hope you have your question answered and that you have a good day.
Ofc, thank you, you too!

NicktheBlade16:

thx